OpenStudy (anonymous):

Whats the product of (7x − 3)(7x + 3).

OpenStudy (diyadiya):

\[\Huge (a-b)(a+b)=a^2-b^2\]

OpenStudy (diyadiya):

\[\Large (7x-3)(7x+3)= \ \ ?\]

OpenStudy (diyadiya):

@Aaron0620

OpenStudy (anonymous):

49x2 - 9 2?

OpenStudy (anonymous):

49x2 + 9?

OpenStudy (diyadiya):

Yes!\[\Large (7x)^2-(3)^2= 49x^2-9\]

OpenStudy (anonymous):

thank you :)

OpenStudy (diyadiya):

welcome!!